A random sample of size n=73 is taken from a population of size n=749 with a population proportion p=0.59 n = 73, p = 0.59 a-1. Is it necessary to apply the finite population correction factor? No a-2. calculate the 1.expected value and the 2.standard error of the sampling proportion
Answer #1

Solution

Given that,

p = 0.59

1 - p = 1 - 0.59 = 0.41

n = 73

1) expected value =   = p = 0.59

2) standard error = =  [p ( 1 - p ) / n] = [(0.59 * 0.41) / 73 ] = 0.0576